Solve the differential equation

(yz+z²) dx - xz dy + xy dz = 0
1
Expert's answer
2019-08-26T09:00:45-0400

Here

"P= yz+z^2 ,Q= -xz ,R= xy""{dx \over yz+z^2}={dy \over -xz}={dz \over xy}"

From 2nd and 3rd fraction, we get 


"{dy \over -z}={dz \over y}""d(y^2+z^2)=0"

Integrating, we get 


"y^2+z^2=c_1"

If we take


"P_1=x, Q_1=z,R_1=-z"

then


"PP_1+QQ_1+RR_1=xyz+xz^2-xz^2-xyz=0"

Thus for the given system of equations, we have 


"{dx \over yz+z^2}={dy \over -xz}={dz \over xy}={xdx+zdy-zdz \over 0}"

"{dx \over x}={dy \over z}={dz \over -z}"

From 1st and 2nd fraction, we get 


"{dx \over x}={dz \over -z}""d(\ln|x|+\ln|z|)=0"

Integrating, we get 


"\ln(|x||z|)=\ln c^*""|x||z|=c_2"

Let F be an arbitrary differentiable function, then the solution of the partial differential equation is


"F(y^2+z^2,|x||z|)=0"